OpenStudy (anonymous):

5.2,8.3,8.5,7.7,7.8,y;mean=7.1

OpenStudy (anonymous):

Find the sum of the elements: 5.2+8.3+8.5+7.7+7.8 = ? What is the number of elements above = ? Mean = (sum of the elements)/(# of elements) Can you do it from here?

OpenStudy (anonymous):

oops misread the question.

OpenStudy (anonymous):

Find the sum of the elements: 5.2+8.3+8.5+7.7+7.8+y = ? What is the number of elements above = ? Mean = 7.1 Mean = (sum of the elements)/(# of elements) Write an equation for the mean from here. Can you do that?

OpenStudy (anonymous):

@ASHLEY_12

OpenStudy (anonymous):

Are u there?

OpenStudy (anonymous):

no i do not no math at all sorry

OpenStudy (anonymous):

Ok can you at least find the answer to the questions below: Find the sum of the elements: 5.2+8.3+8.5+7.7+7.8+y = ? What is the number of elements above = ?

OpenStudy (anonymous):

Sum of elements = 37.5+y # of elements = 6 (5.2, 8.3, 8.5, 7.7, 7.8 and y) Equation: \[7.1 = \frac{ 37.5+y }{ 6 }\] Solve for 'y'